how do you find them? im a bit lost.
Soo Table 21.1 Reaction Mixtures at Room Temperature (Reagent Valumes in ml) Reaction Flask I 250 mL) Reaction Flak Il (125 m this is what we used
A. Dependence of Reaction Rate on Concentration Reaction: 61aq) + BrO,(aq) + 6H(aq)-31,(aq) + Brtaq) +3H,OD ABrO ABro,THTr rate In all the reaction mixtures used in this experiment, the color change occurred when a constant predetermined number of moles of BrO," had been used up in the reaction. The color "clock" allows you to measure the time required for this fixed number of moles of BrO, to react. The rate of cach reaction is determined by the time r required for the color to change; since in Equation 2 the change in concentration of BrO, ion, A[BrO,"1. is the same in cach mixture, the relative rate of cach reaction is inverscly proportional to the time 1. Since we are mainly concerned with relative rather than absolute rates, we willl for convenience take all relative rates as being equal to 1000h. Fill in the following table, first calculating the relative reaction rate for cach mixture. Reactant Concentrations Relative Rate Timer (sec) in Reacting Mixture (M) Temp. Reaction for Color of Reaction H' in ("C Mixture BrO, to Change 1000 4.854 9.615 9.524 ZO6 0.0020 104 2 105 49 20.41 4.545 220 5 The reactant concentrations in the reaction mixture are not those of the stock solutions, since the reagents were diluted by the other solutions. The final volume of the reaction mixnure is 50 ml. in all cases. Since the number of moles of reactant does not change on dilution we can say. for example, for ion, that moles of r IIns X x V For Reaction Mixture 1, M 0.010 M.V10 ml V 50 ml. Therefore, o010 Mx10 ml 50 ml 020 M Calculate the rest of the concentrations in the table using the same approach Determination of the Orders of the Reaction Given the data in the table, the problemi tion. Since we are dealing with relative rates, we can modify Equation 2 to read as follows: is to find the order for each reactant and the rate constant for the reac- relative rate = A' BrO,T[ H' (5) (continued on following page)
Soo Table 21.1 Reaction Mixtures at Room Temperature (Reagent Valumes in ml) Reaction Flask I 250 mL) Reaction Flak Il (125 ml.) Reaction Misture .10 M HC a010 M KI 9.0010 M Na80 HO 0.040 M KBrO 10 10 10 10 10 10 10 20 10 20 10 10 10 10 20 10 10 12 15
Answer #1

Reaction mixture-1:

[BrO3-](mixture) :

Since KBrO3 is completely dissociated, [BrO3-](mixture) = [KBrO3]

V(mixture) = (Volume of all reactants in reaction flask-1) + (Volume of all reactants in reaction flask-2)

= 10+10+10+10+10 = 50 mL (Same for all reaction mixture)

V(stock) = 10 mL

[BrO3-](stock) = 0.010 M (Please confirm whether it is 0.010 M or 0.040 M as the 2nd figure is unclear)

Hence [BrO3-](mixture) = (0.010 M * 10 mL] / 50 mL = 0.0020 M

[H+](mixture):

Since HCl is completely dissociated, [H+](mixture) = [HCl](mixture)

V(mixture) = 50 mL

V(stock) = 10 mL

[H+](stock) = 0.10 M

=> [H+](mixture) = (0.10 M * 10 mL] / 50 mL = 0.020 M

Similarly we can calculate for rest of the reaction mixture as:

Reaction mixture-2:

[I-](mixture) = (0.010 M * 20 mL] / 50 mL = 0.0040 M

[BrO3-](mixture) = (0.010 M * 10 mL] / 50 mL = 0.0020 M

[H+](mixture) = (0.10 M * 10 mL] / 50 mL = 0.020 M

Reaction mixture-3:

[I-](mixture) = (0.010 M * 10 mL] / 50 mL = 0.0020 M

[BrO3-](mixture) = (0.010 M * 20 mL] / 50 mL = 0.0040 M

[H+](mixture) = (0.10 M * 10 mL] / 50 mL = 0.020 M

Reaction mixture-4:

[I-](mixture) = (0.010 M * 10 mL] / 50 mL = 0.0020 M

[BrO3-](mixture) = (0.010 M * 10 mL] / 50 mL = 0.0020 M

[H+](mixture) = (0.10 M * 20 mL] / 50 mL = 0.040 M

Reaction mixture-5:

[I-](mixture) = (0.010 M * 8 mL] / 50 mL = 0.0016 M

[BrO3-](mixture) = (0.010 M * 5 mL] / 50 mL = 0.0010 M

[H+](mixture) = (0.10 M * 15 mL] / 50 mL = 0.030 M

Reactant Concentrations Relative Rate of Reaction Time r (sec) in Reacting Mixture (M) Temp in C) for Color Reaction LHI 0.02